Mercury is consists of 0.146% 197 Hg (195.96583 amu); 10.02% 198 Hg (197.9668 amu); 16.84% 199 Hg (198.9683 amu); 23.13% 200 Hg (199.9683 amu); 13.22% 201 Hg (200.9703 amu); 29.8% 202 Hg (201.97060 amu). Calculate the atomic weight of Hg.


1
Expert's answer
2020-08-29T14:01:48-0400

0,146×197/100=0,29

10,02×198/100=19,84

16,84×199/100=33,51

23,13×200/100=46,26

13,22×201/100=26,57

29,8×202/100=60,2

Ar(Hg)=0,29+19,84+33,51+46,26+26,57+60,2=186,7.